OpenStudy (anonymous):

what the equation of a line in slope intercept form that runs through the points (4,5) and (-2,7)

OpenStudy (jdoe0001):

what would be the slope of such a line?

OpenStudy (lncognlto):

m = y2-y1/x2-x1 = (7-5)/(-2-4) = 2/-6 = -1/3 Subst m = -1/3 and (4,5) in y = mx +c: 5 = -1/3(4) + c therefore c = 19/3 y = -1/3x + 19/3 or 3y = -x + 19